OpenStudy (anonymous):

A set of data items is normally distributed with a mean of 600 and a standard deviation of 20. Find the data item in this distribution that corresponds to the following z-score: z = 1.5

OpenStudy (zarkon):

solve \(\displaystyle \frac{x-\mu}{\sigma}=1.5\) for x

OpenStudy (anonymous):

i got -29.95but i dont think thats right

OpenStudy (zarkon):

\(1.5\times 20+600\)